QUESTION 2 Find the range for the given sample data. Jeanne is currently taking college economics. The instructor often gives quizzes. On the past five quizzes, Jeanne got the following scores: 6 20 1 15 10 Compute the range. O 1 020 O 19 1 p QUESTION 3 Find the percentile for the data point. Data set: 2 3 7425 10 9 11 3 3 10 6 1 10; data point 7 0 70 35
Answer #1

Solution :

Given that,

2 ) Range

= maximum(xi) - minimum(xi)

= 20 - 1

= 19

Range = 19

Answer 19 is correct.

3 )Percentile
Arrange the data in ascending order:
1, 2, 2, 3, 3, 3, 4, 5, 6, 7, 9, 10, 10, 10, 11
Compute the position of the pth percentile (index i):
i = (p / 100) * n),
where p = 7 and n = 15
i = (7 / 100) * 15
= 1.05
The 7th percentile is the value in 2th position, or 2
Let's say we want to determine the 80th percentile for the following values:
10, 55, 50, 80, 85, 90, 20, 40, 55, 65, 30, 25.

Arrange the values from lowest to highest: 10, 20, 25, 30, 40, 50, 55, 55, 65, 80, 85, 90.

Calculate index i (the position of the pth percentile):

i = (p / 100) * n = 80 / 100 * 12 = 9.6

In this case, i is not an integer. As such, you should round up. The point of the 80th percentile is the next integer above 9.6, which is the 10th position.
The 80th percentile is the data value that appears in the 10th position in the list of values: 80.

Let's look at a second example. This time, we want to calculate the 50th percentile for the set of data.

Having arranged the values in ascending order, we have: 10, 55, 50, 80, 85, 90, 20, 40, 55, 65, 30, 25.

Calculate index i (the position of the pth percentile):

i = (p / 100) * n = 50 / 100 * 12 = 6

In this case, i is an integer. As such, the 50th percentile is the average of the values that appear in the 6th and 7th positions in the set of ordered data: 50 and 55 respectively.

Therefore, the 50th percentile of this data set is (50 + 55) / 2 = 52.5.

Answer 52 to is correct.